Solve the following using the method of undetermined coefficients. (d ^ 2 * y)/(d * t ^ 2) + (dy)/(dt) - 5y = - 6x ^ 3 + 3x ^ 2 + 6x


1
Expert's answer
2022-04-21T13:59:38-0400

"y''+y'-5y=-6x^3+3x^2+6x" (1)

The solutions to a nonhomogeneous equation are of the form

"y(x)= y_h(x) + y_p(x)" ,

where "y_h" is the general solution to the associated homogeneous equation and "y_p" is a particular solution.

The associated homogeneous equation:

"y''+y'-5y=0"

The general solution of this equation is determined by the roots of the characteristic equation:

"\lambda^2+\lambda-5=0"

"D=1^2-4\cdot1\cdot(-5)=21"

"\lambda=\frac{-1\pm\sqrt{21}}{2}"

"y_h(x)=Ae^{\frac{-1+\sqrt{21}}{2}x}+Be^{\frac{-1-\sqrt{21}}{2}x}" , where "A" and "B" are constants.

The particular solution of the differential equation:

"y_p(x)=Cx^3+Dx^2+Ex+F"

"y_p'(x)=3Cx^2+2Dx+E"

"y_p''(x)=6Cx+2D"

Substitution "y_p" , "y_p'" , and "y_p''" into (1) gives:

"6Cx+2D+3Cx^2+2Dx+E-""5(Cx^3+Dx^2+Ex+F)=-6x^3+3x^2+6x"

The last equation must be valid for all values of x, so the coefficients with the like powers of x

 in the right and left sides must be identical:

"\begin{cases}
 -5C=-6\\
 3C-5D=3\\
 6C+2D-5E=6\\
 2D+E-5F=0
\end{cases}"

We find from this system that 

 "C=\frac65" ; "D=\frac{3}{25}" ; "E=\frac{36}{125}" ; "F=\frac{66}{625}"

 As a result, the particular solution is written as

"y_p(x)=\frac65x^3+\frac{3}{25}x^2+\frac{36}{125}x+\frac{66}{625}"

Answer: "y(x)=Ae^{\frac{-1+\sqrt{21}}{2}x}+Be^{\frac{-1-\sqrt{21}}{2}x}+\frac65x^3+""\frac{3}{25}x^2+\frac{36}{125}x+\frac{66}{625}".